OpenStudy (anonymous):

. Below are the steps to solve an equation: Step 1: |x - 4| + 1 = 8 Step 2: |x - 4| = 8 - 1 Step 3: |x - 4| = 7 Which of the following is a correct next step to solve the equation? (4 points) x - 4 = -7 x + 4 = -7 -x - 4 = 7 x + 4 = 7

OpenStudy (jdoe0001):

\(\bf |x - 4| = 7\implies \begin{cases} +(x - 4) = 7\implies x-4=7\\ \quad \\ \bf -(x - 4) = 7\implies x-4=-7 \end{cases}\)

OpenStudy (jdoe0001):

absolute value equations have 2 cases or scenarios, and thus "x" will have 2 values, one for each scenario